OpenStudy (anonymous):

solve equation for x logx+log(x-2)=logx3

OpenStudy (mertsj):

Remember that log(a) + log(b)=log(ab) Use that property on the left side.

OpenStudy (anonymous):

what does it mean logx3? cube?

OpenStudy (anonymous):

log3x*

OpenStudy (mertsj):

\[\log(x)^3\]

OpenStudy (mertsj):

Does it look like that?

OpenStudy (mertsj):

Rini. did you do what I said to the left side?

OpenStudy (anonymous):

x (x -2) = 3x x^2 - 2x - 3x = 0 x^2 - 5x = 0 => x = 0, x = 5